OpenStudy (anonymous):

Given a genetic recessive disorder which occurs in 4 of 10,000 individuals, you are asked to find the allele, phenotype and genotype frequencies. Show your work and equations and provide answers in full sentences. Note: use ^2 to indicate squares and sqrt() to indicate taking the square root. (Points : 6)

OpenStudy (anonymous):

@homeSchooler000

OpenStudy (anonymous):

Could you help explain this one as well

OpenStudy (aaronq):

from the Hardy-Weinberg equation (below) and the information given, we find: \(\Large (p+q)^2=q^2+2pq+q^2=1\) -------------------------------------- p= frequency of dominant allele q= frequency of recessive allele \(q^2\)= dominant individual frequency 2pq = heterozygous individual frequency \(p^2\)= recessive individual frequency -------------------------------------- \(q^2\)= frequency of affected individual = \(\dfrac{4}{10000}=0.0004\) q = allele frequency \(q^2=0.0004\rightarrow q=\sqrt{0.0004}=0.02\) Thus the recessive allele frequency is 0.02 and the phenotype frequency is 0.0004, the genotype frequency is the same as the phenotype frequency in this case. Use the equation above to find the rest of the information required.